A wire carrying a steady (DC) 30-A current has length L =12 cm between the pole faces of a magnet. The wire is at angle ϴ=600 to the field. The magnetic field is approximately uniform at 0.90 T. We ignore the field beyond the pole pieces. Determine the magnitude and direction of the magnetic force on the wire. 


1
Expert's answer
2021-07-02T09:33:20-0400

force on a current carrying conductor, "F=IlBSin\theta=30\times 12/100\times 0.9\times Sin60^0=2.80N"